OpenStudy (anonymous):

a construction explosion has an intensity I of 1.25*10^-4W/m. find the loudness of the sound in decibels if L=10 log I/Izero and Izero=10^-12W/m^2 (@zepdrix)

zepdrix (zepdrix):

Hmm it didn't work :) prolly cause you put it in ( ) lol XD I was curious, for the last problem, which method made more sense? Was doing it with brackets easier?

OpenStudy (anonymous):

yes the brackets made it easier

zepdrix (zepdrix):

\[\large I=1.25 \times 10^{-4} \; W/m\] \[\large L=10 \log \left(\frac{I}{I_o}\right)\] \[\large I_o=10^{-12}\;W/m^2\] Did I enter the log correctly? It's just a little bit hard to read in text.

zepdrix (zepdrix):

So this one will be even more painful on your calculator lol :)

zepdrix (zepdrix):

10, log(, (, 1.25, x, 10, ^, (-), 4, ), ÷, (, 10, ^, (-), 12, ), ), I hope the commas aren't making it super confusing. I'm trying to separate the key presses.

zepdrix (zepdrix):

See the little (-) symbol? It's near your enter key. To apply a negative to a number, you CANNOT use the SUBTRACTION key, you have to use the (-) key.

zepdrix (zepdrix):

Hmm I got 80.97. Hmmm let's see :o

zepdrix (zepdrix):

When you get it all punched in, it should look like,\[\huge 10\log((1.25 \times 10\text{^}-4)\div(10\text{^}-12))\]

zepdrix (zepdrix):

Make sure you're not missing the extra bracket at the start of your log.

zepdrix (zepdrix):

Yah if you put brackets around the -4, then it might be multiplying the 1.25 x 10 BEFORE doing the exponent for some reason. Which is backwards.

zepdrix (zepdrix):

lol :) I really wish teachers would put more emphasis on how to use a calculator. Just a few class periods maybe.. I see so many people who struggle with those darn things D:

OpenStudy (anonymous):

I surely do and Im home schooled and the teacher this year just doesn't explain more than what the book says.. I need examples or something to follow. My teachers in the past were great about doing several problems to help.